# Using Reinforcement Learning to Train Large Language Models to Explain Human Decisions

### v1

@@ Abstract
A central goal of cognitive modeling is to develop models that not only predict human behavior but also provide insight into the underlying cognitive mechanisms. While neural network models trained on large-scale behavioral data often achieve strong predictive performance, they typically fall short in offering interpretable explanations of the cognitive processes they capture. In this work, we explore the potential of pretrained large language models (LLMs) to serve as dual-purpose cognitive models—capable of both accurate prediction and interpretable explanation in natural language. Specifically, we employ reinforcement learning with outcome-based rewards to guide LLMs toward generating explicit reasoning traces for explaining human risky choices. Our findings demonstrate that this approach produces high-quality explanations alongside strong quantitative predictions of human decisions.
@@ 1 Introduction
Computational models of cognition have driven immense progress in understanding the mental processes underlying learning, thinking, problem solving, and decision making farrell2018computational ; busemeyer2010cognitive ; griffiths2024bayesian ; ma2023bayesian ; rumelhart1986general . Recent advances–particularly those leveraging neural networks to predict human behavior–have introduced increasingly sophisticated model architectures and training methods peterson2021using ; zhu2024capturing ; huang2023quasi ; almaatouq2024beyond . While these models have demonstrated improved predictive accuracy, they often lack interpretability, offering limited insight into the underlying cognitive mechanisms. A recent example is the Centaur model, which applies customized supervised fine-tuning (SFT) to pretrained LLMs and achieves impressive performance in predicting human behavior, surpassing domain-specific cognitive models from the literature binz2024centaur . However, despite their predictive successes, such models offer limited explanatory power; a deeper theoretical understanding of human cognition requires more than a simple match in behavior frank2025cognitive .
@@
In this work, we put these ideas to the test in the domain of human risky choice. That is, how people choose between risky and/or safe options. To encourage LLMs to generate both CoTs and predictions about human risky choices, we post-train a backbone LLM using reinforcement learning (RL) with outcome-based rewards lambert2025reinforcement ; shao2024deepseekmath ; yu2025dapo ; liu2025understanding , where human choice proportions served as the ground truth for evaluation. To contextualize the RL approach, we also evaluated various SFT post-training methods, including Centaur-style SFT binz2024centaur .
@@ 2 Related Work
Leveraging LLMs to model human cognition. In recent years, there has been widespread enthusiasm about the potential of LLMs to advance cognitive modeling and provide new theoretical understanding about the mind. A growing body of work has sparked scholarly debate around the use of LLMs for modeling human cognitive processes frank2025cognitive ; griffiths2024bayesian ; binz2023turning . The main advantage of using LLMs to predict human behavior is that LLMs can process similar stimuli to people; in other words, LLMs can process a broader range of stimuli (often described in natural language) than previous neural network models which typically operate in a more abstract representation of cognitive tasks.
Consider the risky choice problem illustrated in Figure 1 as an example. Traditional cognitive models and neural-network-based models typically operate on structured quantitative task features such as $\{1,\$27\}$ (i.e., probability and value) for option A and $\{0.9,\$25;0.1,\$92\}$ for option B peterson2021using . These models define functions that map from such numerical inputs to human risky choices. In contrast, LLM-based cognitive models operate over a different representation of the task: they take natural language descriptions as input and predict human responses directly zhu2024language ; binz2024centaur ; binz2023turning . Recent work has shown that SFT of LLMs can improve predictive accuracy on human behavioral data binz2024centaur .
Automated discovery of cognitive models using LLMs. Another closely-related line of research involves using LLMs to automatically search over the space of cognitive models, enabling the automated discovery of interpretable theories castro2025discovering ; rmus2025towards . In this approach, LLMs are prompted to generate symbolic programs (i.e., Python code), which are then executed and fitted to human or animal data. Because the discovered models are expressed as code, they are inherently interpretable–helping to address the aforementioned dilemma in cognitive modeling. Leveraging the coding capabilities of LLMs has shown promise in identifying heuristic decision-making models rmus2025towards and strategies in multi-armed bandit tasks castro2025discovering . However, these approaches typically do not involve further fine-tuning of the LLM during the search process, instead relying heavily on the model’s in-context learning brown2020language ability for effective model discovery.
@@ 3 Method
To evaluate whether post-training LLMs can produce useful cognitive models, we compare three distinct post-training strategies for LLMs: (i) SFT, (ii) a variant of SFT specifically designed for adapting LLMs to cognitive tasks, as used in the Centaur model binz2024centaur , and (iii) RL based on Group Relative Policy Optimization (GRPO) shao2024deepseekmath ; liu2025understanding . Each method was applied to fine-tune identical low-rank adaptation (LoRA) modules on the largest available human risky-choice dataset, choices13k, originally collected by peterson2021using .
Figure 1: Overview of three post-training strategies for building cognitive models of human risky choice using Qwen-2.5-7B-Instruct. The backbone LLM was first adapted using low-rank adaptation (LoRA) hu2022lora , followed by post-training via three strategies: supervised fine-tuning (SFT), Centaur-style SFT binz2024centaur , and reinforcement learning from outcome-based rewards liu2025understanding ; shao2024deepseekmath . In the illustrated example, the LLM is prompted to predict human choice behavior. SFT and Centaur-style models are trained directly on human data represented in JSON format. In contrast, the RL model generates candidate completions that include CoT reasoning and final predictions in JSON format, with each completion evaluated based on its predictions. Tokens or predictions relevant to each training method are highlighted in light blue.
The backbone LLM used throughout is Qwen-2.5-7B-Instruct yang2024qwen2 . All models were fine-tuned using the same LoRA configuration hu2022lora , with both the rank and alpha parameters set to 32. LoRA modules were applied to all linear layers of the backbone model and included a dropout probability of 0.05. This setup results in approximately 80.74M trainable parameters, comprising only 1.05% of the total parameters in the 7B model.
@@
We fine-tuned the Qwen model using SFT for a total of 6 epochs on the training set, with a fixed learning rate of $10^{-5}$, gradient accumulation steps of 8, and the AdamW optimizer loshchilov2017decoupled . In our setup, SFT corresponds to the standard autoregressive next-token prediction paradigm.
In parallel, we implemented a variant used to train the Centaur model, which selectively masks out all tokens except those corresponding to human data binz2024centaur . This method has been shown to effectively adapt pretrained LLMs for explaining a wide range of human cognitive tasks binz2024centaur . Unlike standard SFT, Centaur-style SFT places human data (i.e., choice proportions) within special brackets (“<<” and “>>”), and only tokens enclosed within these brackets contribute to the loss during training (see Figure 1). In our implementation, Centaur-style SFT was applied at the problem level rather than the individual participant level, such that the model predicts aggregated human responses for each risky-choice problem.
RL training details. We also fine-tuned the Qwen model using the GRPO algorithm shao2024deepseekmath . At each training step, the model generates 12 candidate completions (i.e., $\mathbf{o}_{1},\mathbf{o}_{2},...,\mathbf{o}_{G}$) for the same risky-choice problem (i.e., $\mathbf{q}$), with each completion capped at a maximum length of 1,024 tokens. The GRPO algorithm evaluates each candidate based on the model-predicted choice probabilities for the two options (see Figure 1 RL). Specifically, we implemented a outcome-based reward function that uses human data as ground truth. The reward is defined as one minus the absolute difference between the model’s predicted probability for option B and the empirical proportion of human participants who chose that option:
@@
In addition, we incorporated a format reward to encourage structurally well-formed completions. This reward depends on the number and position of JSON-formatted outputs in the model response. If the model generates exactly one JSON output, the format reward increases by 0.25. Furthermore, if the first JSON output appears after the 500th character–encouraging reasoning tokens to precede the prediction–the format reward increases by an additional 0.25. The maximum possible format reward is therefore 0.5.
The RL training was conducted for a total of 3 epochs, with a learning rate of $3\times 10^{-6}$ and a cosine learning rate scheduler. Moreover, we omitted the reward normalization by standard deviation as suggested by the GRPO Done Right algorithm liu2025understanding . As a result, the advantage function was defined as the reward of each candidate completion minus the average reward within the group.
@@
where $G=12$ is the group size. We revisit the implications of this design choice, along with other unsuccessful attempts, in the Discussion section.
With the advantage function defined, we train the backbone LLM using the GRPO Done Right algorithm with the following objective function:
@@
where $\mathbf{q}$ denotes a risky-choice problem sampled from the training set $Q$, and $|\mathbf{o}_{i}|$ is the number of tokens in the i-th model completion $\mathbf{o}_{i}$, generated by the LLM under the old policy $\pi_{\theta_{\text{old}}}$. Moreover, $t$ denotes the position of a token within the sequence of a model-generated completion and $\theta$ represents the parameters of LLM. The inner summation iterates over all tokens in the sampled completion. This objective follows the clipped surrogate loss formulation from the proximal policy optimization (PPO) algorithm schulman2017proximal , modified to operate at the token level within each sampled trajectory. The clipped values lie within the range of [0.8, 1.28], where $\epsilon_{\text{low}}=0.2$ and $\epsilon_{\text{high}}=0.28$, as set in our experiment. This asymmetric clipping follows the recommendation in yu2025dapo , which suggests that slightly increasing $\epsilon_{\text{high}}$ can enhance exploration in RL. GRPO also incorporates a KL divergence penalty term in its objective function to prevent the updated policy from deviating too far from the backbone LLM. This penalty takes the form $\beta D_{KL}(\pi_{\theta}\parallel\pi_{\text{reference}})$, where $\beta$ is set to $10^{-4}$.
Key distinctions between RL and Centaur-style SFT. While both Centaur-style SFT and RL focus exclusively on tokens relevant to choice probabilities (see Figure 1, highlighted text), there are important distinctions between these two post-training methods. Centaur-style SFT operates within the standard next-token prediction framework and thus relies on the tokenized representation of numerical outputs. In contrast, RL assigns outcome rewards based on the predicted numerical values themselves, rather than their tokenized forms. These rewards are then used to weight policy updates during training. This reward-weighted policy optimization has been argued to support improved generalization in downstream tasks chu2025sft ; wang2025reinforcement .
@@ 4 Results
Having introduced the three post-training strategies, we now evaluate the effectiveness of the fine-tuned models from each method as cognitive models of human risky choice. To ensure comparability, all fine-tuned models were evaluated using identical sampling parameters, implemented with vLLM kwon2023efficient : temperature = 0.7, top-p = 0.95, and top-k = 50. The maximum number of generated tokens was adjusted according to model type. Since the RL models were trained to generate intermediate reasoning before producing a final choice prediction, they were allowed up to 1,024 tokens to accommodate more elaborate completions. In contrast, the SFT and Centaur-style models were restricted to 30 tokens, as they were trained to produce choice predictions directly without intermediate reasoning steps. We evaluated model inferences at all checkpoints across the three post-training methods, for both the training and test set problems.
@@
Figure 2: Learning curves on the (a) training and (b) test sets. Backbone LLM is Qwen-2.5-7B-Instruct. The horizontal axes indicate training epochs, while the vertical axes represent mean squared error (MSE) evaluated on the corresponding dataset. The three post-training strategies compared are supervised fine-tuning (SFT, red), Centaur-style SFT (green), and reinforcement learning (RL, grey). The lowest MSEs on the training set occur at epochs 5.98 (SFT), 5.74 (Centaur-style SFT), and 2.75 (RL), while the lowest MSEs on the test set occur at epochs 5.86 (SFT), 5.86 (Centaur-style SFT), and 2.60 (RL). (c) MSE on the test set at the final checkpoint of each post-training method. Error bars represent ±$\pm 1$ standard error across risky-choice problems.
[…]
@@
Figure 5: Visualization of CoT reasoning generated by RL models. Each reasoning segment (“thought”) is embedded using the all-MiniLM-L6-v2 model from SBERT reimers-2019-sentence-bert , followed by dimensionality reduction to two dimensions using t-SNE van2008visualizing . In the resulting 2D space, we identified nine clusters using the k-means algorithm. Each cluster is labeled with a summary of its centroid thought to provide an interpretable overview of common reasoning patterns.
We next aim to characterize the CoT reasoning learned by the RL model in order to identify effective cognitive processes that contribute to accurately predicting human risky choices. Note that the CoT generated by the RL model is typically itemized, often presented as a sequence of clearly delineated steps or sections. Leveraging this structure, we segmented each CoT into smaller units–referred to as thoughts–using a regular expression-based parser. For example, the CoT shown in Figure 5 (right) is divided into five distinct thoughts, corresponding to its five itemized sections.
Each segmented thought was embedded using the all-MiniLM-L6-v2 model from SBERT reimers-2019-sentence-bert , and subsequently projected into two dimensions using t-SNE for visualization van2008visualizing . In this 2D embedding space, we identified nine clusters using the k-means algorithm. To enhance interpretability, each cluster was labeled with a summary of its centroid thought, providing an overview of common reasoning patterns.
@@
Moreover, we conducted several supplementary analyses of CoT reasoning, including small-scale human evaluations (Appendix C.1), large-scale LLM-as-a-judge assessments (Appendix C.2), and CoT swapping between the backbone and RL models (Appendix C.3). Taken together, these analyses suggest that RL post-training improves the quality of CoT reasoning, which in turn contributes to enhanced prediction accuracy of human behavior.
Cognitive mechanisms identified in CoT. We examined the cognitive mechanisms reflected in the RL model’s CoT reasoning. To assist with this analysis, we used GPT-4.1111https://openai.com/index/gpt-4-1/ to summarize the CoT outputs generated by the RL model (see Appendix B for detailed prompts). The evolution of the top eight cognitive mechanisms over training epochs is shown in Figure 6a. Among these, expected value computation and risk aversion emerged as the two most frequently used mechanisms, each accounting for approximately 29% to 36% of thoughts across risky-choice problems.
We observed notable representations of loss aversion–the tendency for individuals to prefer avoiding losses over acquiring equivalent gains tversky1992advances –and the certainty effect, where individuals disproportionately favor certain outcomes over those that are merely probable cohen1988certainty . Finally, the RL model occasionally considers the possibility that some individuals may exhibit risk-seeking behavior or be influenced by other cognitive biases, such as probability weighting prelec1998probability ; kahneman2013prospect , framing effects tversky1981framing , and ambiguity aversion fox1995ambiguity . The full list of cognitive mechanisms identified by the RL model is provided in Table 1 and Appendix B.
These results also provide an interesting perspective for theorists working on models of human risky choice, a domain where substantial research effort has been devoted to identifying systematic deviations from rationality and developing heuristics and biases to account for human irrational behaviors tversky1992advances ; kahneman2013prospect ; gigerenzer2011heuristic ; gilovich2002heuristics . In contrast, the RL model highlights calculating expected values and risk aversion as the dominant forces driving the explanation and prediction of human risky choices. This finding suggests that greater attention should be paid to rational characterizations of human behavior and that there may be value in developing new theories grounded in rational principles. Recent work in this area suggests that human decisions may be explained in terms of the rational use of cognitive resources gershman2015computational ; lieder2020resource ; zhucomputation .

### v2

@@ Abstract
A central goal of cognitive modeling is to develop models that not only predict human behavior but also provide insight into the underlying cognitive mechanisms. While neural network models trained on large-scale behavioral data often achieve strong predictive performance, they typically fall short in offering interpretable explanations of the cognitive processes they capture. In this work, we explore the potential of pretrained large language models (LLMs) to serve as dual-purpose cognitive models — capable of both accurate prediction and interpretable explanation in natural language. Specifically, we employ reinforcement learning with outcome-based rewards to guide LLMs toward generating explicit reasoning traces that explain human risky choices. Our findings demonstrate that this approach produces high-quality explanations at scale alongside strong quantitative predictions of human decisions.
@@ 1 Introduction
Computational models of cognition have driven immense progress in understanding the mental processes underlying learning, thinking, problem solving, and decision making (Farrell and Lewandowsky, 2018; Busemeyer and Diederich, 2010; Griffiths et al., 2024a; Ma et al., 2023; Rumelhart et al., 1986). Recent advances — particularly those leveraging neural networks to predict human behavior — have introduced increasingly sophisticated model architectures and training methods (Peterson et al., 2021; Zhu et al., 2025a; Huang, 2023; Almaatouq et al., 2024). While these models have demonstrated improved predictive accuracy, they often lack interpretability, offering limited insight into the underlying cognitive mechanisms. A recent example is the Centaur model, which applies customized supervised fine-tuning (SFT) to pretrained LLMs and achieves impressive performance in predicting human behavior, surpassing domain-specific cognitive models from the literature (Binz et al., 2025). However, despite their predictive successes, such models offer limited explanatory power; a deeper theoretical understanding of human cognition requires more than a simple match in behavior (Frank and Goodman, 2025).
@@
In this work, we put these ideas to the test in the domain of human risky choice; that is, how people choose between risky and/or safe options. To encourage LLMs to generate both CoTs and predictions about human risky choices, we post-train a backbone LLM using reinforcement learning (RL) with outcome-based rewards (Lambert, 2025; Shao et al., 2024; Yu et al., 2025; Liu et al., 2025), where human choice proportions served as the ground truth for evaluation. In other words, the behavioral predictions are directly linked to the reward function of RL, incentivizing the LLM to produce useful reasoning chains. For comparison, we also evaluated various SFT post-training methods, including Centaur-style SFT (Binz et al., 2025), which has been shown to produce high-performing cognitive models based on LLMs.
@@ 2 Related Work
Leveraging LLMs to model human cognition. In recent years, there has been widespread enthusiasm about the potential of LLMs to advance cognitive modeling and provide new theoretical understanding about the mind. A growing body of work has sparked scholarly debate around the use of LLMs for modeling human cognitive processes (Frank and Goodman, 2025; Griffiths et al., 2024b; Binz and Schulz, 2023). The main advantage of using LLMs to predict human behavior is that LLMs can process similar stimuli to people; in other words, LLMs can process a broader range of stimuli (often described in natural language) than previous neural network models which typically operate in a more abstract representation of cognitive tasks (Frank and Goodman, 2025; Peterson et al., 2021; Huang, 2023; Zhu et al., 2025a).
Consider the risky-choice problem illustrated in Figure 1, where human participants were asked to choose between Option A, which offers \$27 for sure, and Option B, which offers \$25 with a 90% probability and \$92 with a 10% probability. Traditional cognitive models and neural-network-based models typically operate on structured quantitative task features such as $\{1,\mathdollar 27\}$ (i.e., probability and value) for option A and $\{0.9,\mathdollar 25;0.1,\mathdollar 92\}$ for option B (Peterson et al., 2021). These models define functions that map from such numerical inputs to human risky choices. In contrast, LLM-based cognitive models operate over a different representation of the task: they take natural language descriptions as input and predict human responses directly (Zhu et al., 2025b; Binz et al., 2025; Binz and Schulz, 2023). Recent work has shown that SFT of LLMs can improve predictive accuracy on human behavioral data (Binz et al., 2025). Our work goes beyond improving LLMs’ ability to predict human behavior; we aim to elicit verbal theories of human behavior from LLMs using RL.
Automated discovery of cognitive models using LLMs. Another closely-related line of research involves using LLMs to automatically search over the space of cognitive models, enabling the automated discovery of interpretable theories (Castro et al., 2025; Rmus et al., 2025; Musslick et al., 2024; Wong et al., 2023). In this approach, LLMs are prompted to generate symbolic programs (e.g., Python code), which are then executed and fitted to human or animal data. Because the discovered models are expressed as code, they are inherently interpretable — helping to address the aforementioned dilemma in cognitive modeling. Leveraging the coding capabilities of LLMs has shown promise in identifying heuristic decision-making models (Rmus et al., 2025) and strategies in multi-armed bandit tasks (Castro et al., 2025). However, these approaches typically do not involve further fine-tuning of the LLM during the search process, instead relying heavily on the model’s in-context learning (Brown et al., 2020) ability for effective model discovery.
@@ 3 Method
To evaluate whether post-training LLMs can produce useful cognitive models, we compare three distinct post-training strategies for LLMs: (i) SFT, (ii) a variant of SFT specifically designed for adapting LLMs to cognitive tasks, as used in the Centaur model (Binz et al., 2025), and (iii) RL based on Group Relative Policy Optimization (GRPO) (Shao et al., 2024; Liu et al., 2025). Each method was applied to fine-tune identical low-rank adaptation (LoRA) modules on the largest available human risky-choice dataset, choices13k, originally collected by Peterson et al. (2021).
Figure 1: Overview of three post-training strategies for building cognitive models of human risky choice using Qwen-2.5-7B-Instruct. The backbone LLM was first adapted using low-rank adaptation (LoRA) (Hu et al., 2022), followed by post-training via three strategies: supervised fine-tuning (SFT), Centaur-style SFT (Binz et al., 2025), and reinforcement learning from outcome-based rewards (Liu et al., 2025; Shao et al., 2024). In the illustrated example, the LLM is prompted to predict human choice behavior. SFT and Centaur-style models are trained directly on human data represented in JSON format. In contrast, the RL model generates candidate completions that include CoT reasoning and final predictions in JSON format, with each completion evaluated based on its predictions. Tokens or predictions relevant to each training method are highlighted in light blue.
The backbone LLM used throughout is Qwen-2.5-7B-Instruct (Yang et al., 2024). All models were fine-tuned using the same LoRA configuration (Hu et al., 2022), with both the rank and alpha parameters set to 32. LoRA modules were applied to all linear layers of the backbone model and included a dropout probability of 0.05. This setup results in approximately 80.74M trainable parameters, comprising only 1.05% of the total parameters in the 7B model.
@@
We fine-tuned the Qwen model using SFT for a total of 6 epochs on the training set, with a fixed learning rate of $10^{-5}$, gradient accumulation steps of 8, and the AdamW optimizer (Loshchilov and Hutter, 2017). In our setup, SFT corresponds to the standard autoregressive next-token prediction paradigm.
In parallel, we implemented a variant used to train the Centaur model, which selectively masks out all tokens except those corresponding to human data (Binz et al., 2025). This method has been shown to effectively adapt pretrained LLMs for explaining a wide range of human cognitive tasks (Binz et al., 2025). Unlike standard SFT, Centaur-style SFT places human data (i.e., choice proportions) within special brackets (“<<<<” and “>>>>”), and only tokens enclosed within these brackets contribute to the loss during training (see Figure 1). In our implementation, Centaur-style SFT was applied at the problem level rather than the individual participant level, such that the model predicts aggregated human responses for each risky-choice problem.
RL training details. We also fine-tuned the Qwen model using the GRPO algorithm (Shao et al., 2024). At each training step, the model generates 12 candidate completions (i.e., $\mathbf{o}_{1},\mathbf{o}_{2},...,\mathbf{o}_{G}$) for the same risky-choice problem (i.e., $\mathbf{q}$), with each completion capped at a maximum length of 1,024 tokens. The GRPO algorithm evaluates each candidate based on the model-predicted choice probabilities for the two options (see Figure 1 RL). Specifically, we implemented an outcome-based reward function that uses human data as ground truth. The reward is defined as one minus the absolute difference between the model’s predicted probability for option B and the empirical proportion of human participants who chose that option:
@@
In addition, we incorporated a format reward to encourage structurally well-formed completions. This reward depends on the number and position of JSON-formatted outputs in the model response. If the model generates exactly one JSON output, the format reward increases by 0.25. Furthermore, if the behavioral prediction appears after reasoning tokens, the format reward increases by an additional 0.25. The maximum possible format reward is therefore 0.5.
The RL training was conducted for a total of 3 epochs, with a learning rate of $3\times 10^{-6}$ and a cosine learning rate scheduler. Moreover, we omitted the reward normalization by standard deviation as suggested by the GRPO Done Right algorithm (Liu et al., 2025). As a result, the advantage function was defined as the reward of each candidate completion minus the average reward within the group.
@@
where $G=12$ is the group size. We revisit the implications of this design choice, along with other unsuccessful attempts, in the Discussion section and Appendix F.
With the advantage function defined, we train the backbone LLM using the GRPO Done Right algorithm (Liu et al., 2025) with the following objective function:
@@
where $\mathbf{q}$ denotes a risky-choice problem sampled from the training set $Q$, and $|\mathbf{o}_{i}|$ is the number of tokens in the i-th model completion $\mathbf{o}_{i}$, generated by the LLM under the old policy $\pi_{\theta_{\text{old}}}$. Moreover, $t$ denotes the position of a token within the sequence of a model-generated completion and $\theta$ represents the parameters of LLM. The inner summation iterates over all tokens in the sampled completion. This objective follows the clipped surrogate loss formulation from the proximal policy optimization (PPO) algorithm (Schulman et al., 2017), modified to operate at the token level within each sampled trajectory. The clipped values lie within the range of [0.8, 1.28], where $\epsilon_{\text{low}}=0.2$ and $\epsilon_{\text{high}}=0.28$, as set in our experiment. This asymmetric clipping follows the recommendation in Yu et al. (2025), which suggests that slightly increasing $\epsilon_{\text{high}}$ can enhance exploration in RL. GRPO also incorporates a KL divergence penalty term in its objective function to prevent the updated policy from deviating too far from the backbone LLM. This penalty takes the form $\beta D_{KL}(\pi_{\theta}\parallel\pi_{\text{reference}})$, where $\beta$ is set to $10^{-4}$.
Key distinctions between RL and Centaur-style SFT. While both Centaur-style SFT and RL focus exclusively on tokens relevant to choice probabilities (see Figure 1, highlighted text), there are important distinctions between these two post-training methods. Centaur-style SFT operates within the standard next-token prediction framework and thus relies on the tokenized representation of numerical outputs. In contrast, RL assigns outcome rewards based on the predicted numerical values themselves, rather than their tokenized forms. These rewards are then used to weight policy updates during training. This reward-weighted policy optimization has been argued to support improved generalization in downstream tasks (Chu et al., 2025; Wang et al., 2025).
@@ 4 Results
Having introduced the three post-training strategies, we now evaluate the effectiveness of the fine-tuned models from each method as cognitive models of human risky choice. To ensure comparability, all fine-tuned models were evaluated using identical sampling parameters, implemented with vLLM (Kwon et al., 2023): temperature = 0.7, top-p = 0.95, and top-k = 50. The maximum number of generated tokens was adjusted according to model type. Since the RL models were trained to generate intermediate reasoning before producing a final choice prediction, they were allowed up to 1,024 tokens to accommodate more elaborate completions. In contrast, the SFT and Centaur-style models were restricted to 30 tokens, as they were trained to produce choice predictions directly without intermediate reasoning steps. We evaluated model inferences at all checkpoints across the three post-training methods, for both the training and test set problems.
@@
Figure 2: Learning curves on the (a) training and (b) test sets. Backbone LLM is Qwen-2.5-7B-Instruct. The horizontal axes indicate training epochs, while the vertical axes represent mean squared error (MSE) evaluated on the corresponding dataset. The three post-training strategies compared are supervised fine-tuning (SFT, red), Centaur-style SFT (green), and reinforcement learning (RL, grey). The lowest MSEs on the training set occur at epochs 5.98 (SFT), 5.74 (Centaur-style SFT), and 2.75 (RL), while the lowest MSEs on the test set occur at epochs 5.86 (SFT), 5.86 (Centaur-style SFT), and 2.60 (RL). (c) MSE on the test set at the final checkpoint of each post-training method. Error bars represent $\pm 1$ standard error across risky-choice problems.
@@
Ultimately, all three post-training methods yield fine-tuned models with comparable performance in predicting human choices. At the final checkpoint, the MSE on test set is M=.0144(SE=.0006)M=.0144~(SE=.0006) for SFT, M=.0155(SE=.0006)M=.0155~(SE=.0006) for Centaur-style SFT, M=.0148(SE=.0006)M=.0148~(SE=.0006) for RL. Statistical tests reveal no significant differences among the models: SFT vs. Centaur-style SFT (t(2923)=1.31,p=0.19t(2923)=-1.31,p=0.19), SFT vs. RL (t(2923)=0.58,p=0.56t(2923)=-0.58,p=0.56), and RL vs. Centaur-style SFT ($t(2923)=0.78,p=0.43$). For reference, the best-performing neural network model reported in (Peterson et al., 2021), the Mixture of Theories model, achieved an MSE of .0113.0113, while the neurally augmented prospect theory model achieved an MSE of .0204.0204 (both models use problem features as input rather than text).
@@
Figure 5: Visualization of CoT reasoning generated by RL models. Each reasoning segment (“thought”) is embedded using the all-MiniLM-L6-v2 model from SBERT (Reimers and Gurevych, 2019), followed by dimensionality reduction to two dimensions using t-SNE (Van der Maaten and Hinton, 2008). In the resulting 2D space, we identified nine clusters using the k-means algorithm. Each cluster is labeled with a summary of its centroid thought to provide an interpretable overview of common reasoning patterns.
We next aim to characterize the CoT reasoning learned by the RL model in order to identify effective cognitive processes that contribute to accurately predicting human risky choices. Note that the CoT generated by the RL model is typically itemized, often presented as a sequence of clearly delineated steps or sections. Leveraging this structure, we segmented each CoT into smaller units — referred to as thoughts — using a regular expression-based parser. For example, the CoT shown in Figure 5 (right) is divided into five distinct thoughts, corresponding to its five itemized sections.
Each segmented thought was embedded using the all-MiniLM-L6-v2 model from SBERT (Reimers and Gurevych, 2019), and subsequently projected into two dimensions using t-SNE for visualization (Van der Maaten and Hinton, 2008). In this 2D embedding space, we identified nine clusters using the k-means algorithm. To enhance interpretability, each cluster was labeled with a summary of its centroid thought, providing an overview of common reasoning patterns.
@@
Moreover, we conducted several supplementary analyses of CoT reasoning, including small-scale human evaluations (Appendix C.1), large-scale LLM-as-a-judge assessments (Appendix C.2), CoT swapping between the backbone and RL models (Appendix C.3), and tests of the effectiveness of SFT-generated and Centaur-generated CoTs (Appendix C.4). Taken together, these analyses suggest that RL post-training improves the quality of CoT reasoning, which in turn contributes to enhanced prediction accuracy of human behavior.
Cognitive mechanisms identified in CoT. We examined the cognitive mechanisms reflected in the RL model’s CoT reasoning. In our context, cognitive mechanisms are specifically defined as verbal theories that explain how people reach a risky decision when presented with a particular set of options. To assist with this analysis, we used OpenAI’s GPT-4.1 (https://openai.com/index/gpt-4-1/) to summarize the CoT outputs generated by the RL model (see Appendix B for detailed prompts). The evolution of the top eight cognitive mechanisms over training epochs is shown in Figure 6a. Among these, expected value computation and risk aversion emerged as the two most frequently used mechanisms, each accounting for approximately 29% to 36% of thoughts across risky-choice problems.
We observed notable representations of loss aversion — the tendency for individuals to prefer avoiding losses over acquiring equivalent gains (Tversky and Kahneman, 1992) — and the certainty effect, where individuals disproportionately favor certain outcomes over those that are merely probable (Cohen and Jaffray, 1988). Finally, the RL model occasionally considers the possibility that some individuals may exhibit risk-seeking behavior or be influenced by other cognitive biases, such as probability weighting (Prelec, 1998; Kahneman and Tversky, 2013), framing effects (Tversky and Kahneman, 1981), and ambiguity aversion (Fox and Tversky, 1995). The full list of cognitive mechanisms identified by the RL model is provided in Table 1 and Appendix B.
These results also provide an interesting perspective for theorists working on models of human risky choice, a domain where substantial research effort has been devoted to identifying systematic deviations from rationality and developing heuristics and biases to account for human irrational behaviors (Tversky and Kahneman, 1992; Kahneman and Tversky, 2013; Gigerenzer and Gaissmaier, 2011; Gilovich et al., 2002). In contrast, the RL model highlights calculating expected values and risk aversion as the dominant forces driving the explanation and prediction of human risky choices. This finding suggests that greater attention should be paid to rational characterizations of human behavior and that there may be value in developing new theories grounded in rational principles. Recent work in this area suggests that human decisions may be explained in terms of the rational use of cognitive resources (Gershman et al., 2015; Lieder and Griffiths, 2020; Zhu and Griffiths, 2025).